OpenStudy (sloppycanada):

Write the standard form equation of a circle with a center at (-5, 5) and a radius of sqrt of 7. For an equation like this, doesn't the radius equal 7? So the equation looks like the following - (x+5)^2 + (y-5)^2 = 7

OpenStudy (phi):

the equation for a circle is \[ (x-h)^2 + (y-k)^2 = r^2 \] where (h,k) is the center, and r is the radius your equation (x+5)^2 + (y-5)^2 = 7 is correct But ***doesn't the radius equal 7?** no, they tell you the radius is \( \sqrt{7}\). However, \( r^2= 7\)

OpenStudy (sloppycanada):

But I simplify it to 7, right?

OpenStudy (phi):

*** But I simplify it to 7, right?*** if you mean \[ r^2= \sqrt{7}\cdot \sqrt{7}= 7 \] then yes. but the radius is \( \sqrt7\) (about 2.646 )